# MOND as a transformation between non-inertial reference frames via Sciama’s interpretation of Mach’s Principle

### v3

@@ 1.2 Mach’s Principle and modified inertia
Sciama’s relationship with today’s measurements of the current cosmological model is satisfied only in terms of orders of magnitude, because the Hubble tension impedes a precise calculation. It yields the orders of magnitude of the observed gravitational constant when considering not the baryonic and dark matter content for $M_{u}$, but the total energy content (since all forms of energy must gravitate according to GR), which roughly corresponds to the critical density of the universe as measured to be almost flat. As Milgrom stated in one of his conferences, “the only system that is strongly general relativistic and in the MOND regime is the Universe at large”. However, the critical density is calculated with the assumption of a constant gravitational constant itself and considering the dark matter content, which could be reduced through an effective correction to the gravitational constant, as will be proposed in the next section. Thus, an honest calculation must be done with direct observations only by the total luminosity of the universe and its average mass-to-light ratio, which yields a value $10^{52}kg<M_{u}<10^{53}kg$, and the observable radius of the universe, which can be estimated by direct observations of the Hubble parameter (subjected to the Hubble tension), which yields around $R_{u}\sim 10^{26}m$.
@@
Alexander Unzicker explored galaxy dynamics in relation to Mach’s Principle and the rotating bucket of water experiment, and proposed a basis for a MOND to solve flat rotation curves without considering Milgrom’s MOND [26]. He proposes the following thought experiment. According to Mach, the system of two distant masses rotating around their center of mass in equilibrium of gravitational and centrifugal forces in an empty universe, since rotation between them is undefined due to absence of absolute space, must be equivalent to the system of those two distant masses not rotating and without gravity (only radial or relative velocities are measurable, and tangential velocities are impossible to measure instantaneously, as they are indistinguishable from a rotation of the coordinate system of the observer). The maximum possible angular velocity is limited by $w_{max}=c/r$ so that the maximum tangential speed is $v=c$ for one of the masses rotating around the other one, since the speed of light cannot be surpassed in any case. This luminal rotation in absence of a background universe from Unzicker’s thought experiment is a feature of introducing Sciama’s relationship in Newton’s law of gravity, in which the Newtonian $v=\sqrt{GM/r}$ changes to $v=c\sqrt{(M/r)/(M_{u}/R_{u})}$ after introducing $G=c^{2}/(M_{u}/R_{u})$ and without a background universe $M_{u}\rightarrow M$ and $R_{u}\rightarrow r$, velocities are $v=c$.
@@
Even though interesting ideas have been put forward relating MOND to Mach’s principle, an exact equivalence and its implications will be presented onwards, which should arise as an approximation in any non-linear theory of modified inertia or gravity which attempts to explain dark matter effects in galaxies. The motivation behind this interpretation of Milgrom’s MOND is that it can be reformulated with the same parameters of the mass and radius of the observable and causally connected universe that the Machian models of modified inertia introduce.
@@ 2 Milgrom’s MOND from Mach’s Principle
Starting from the well-known form of Milgrom’s MOND standard or simple interpolating function correction to Newtonian gravity $g_{N}=GM/r^{2}$ or to inertia $F=m_{i}g$, the ‘true’ gravitational acceleration $g$ or ‘true’ second law of inertia is
$g=\frac{G}{\mu\left(\frac{g_{N}}{a_{0}}\right)}\frac{M}{r^{2}}\ \ or\ \ F=\mu\left(\frac{g}{a_{0}}\right)m_{i}g$ (1)
@@
with $a_{0}$ being Milgrom’s MOND acceleration scale constant, $g_{N}=GM/r^{2}$ the Newtonian gravitational acceleration, $M$ the active gravitational mass, $m_{i}$ the inertial mass, and $r$ the distance between the centers of mass of the active and passive gravitational masses, so that for the motion of a test particle in a gravitational field in the low acceleration or deep MOND regime, $\mu\left(a/a_{0}\right)\rightarrow a/a_{0}$ and $g=\sqrt{a_{0}\ a}$, so that with $g=v^{2}/r$, $v=\sqrt[4]{GMa_{0}}$ for achieving flat galaxy rotation curves in agreement with the Tully-Fisher law in a remarkable simple way with a single new free parameter.
Milgrom’s MOND acceleration scale constant $a_{0}$ is usually interpreted as a fundamental constant in MOND theory, but possibly related to the cosmological parameter of the cosmological constant $a_{0}\sim c^{2}\sqrt{\Lambda}$ or the energy density of the vacuum $\rho_{vac}$ with $\Lambda=8\pi G\rho_{vac}/c^{2}$ according to GR, which coincides in orders of magnitude (a better match can be done with $a_{0}=c^{2}\sqrt{\Lambda/3}$, although integers and Pi are onwards omitted for simplicity). This is due to Milgrom’s wish to define absolute acceleration with respect to the vacuum energy frame in MOND as modified inertia. But Mach’s principle implies that acceleration should be defined with respect to the rest of the masses of the universe. In order to express $a_{0}$ in terms of the mass and radius of the universe, $\rho_{vac}$ is substituted for the matter density of the universe $\rho_{vac}\sim\rho_{u}\sim M_{u}/{R_{u}}^{3}$ with $M_{u}$ the causally connected mass to the local system of study through the radius $R_{u}$ of the observable universe (or particle horizon radius), a relationship which also coincides in orders of magnitude. Newton’s gravitational constant $G$ is substituted for Sciama’s interpretation of Mach’s principle $G\sim c^{2}/(M_{u}/R_{u})$, so that $a_{0}\sim GM_{u}/{R_{u}}^{2}\sim c^{2}/R_{u}$, and by reverse engineering, $v\sim c\sqrt[4]{M/M_{u}}$ and $g\sim c^{2}\sqrt{M/M_{u}}/r$ in the equivalent deep MOND regime. The interpolating function is now scale-invariant in mass and length, resulting in a Machian MOND formulation
@@
$\mu\left(\sqrt{\frac{M/{r}^{2}}{M_{u}/{R_{u}}^{2}}}\right)=\frac{1}{\sqrt{1+\left(\frac{M_{u}/{R_{u}}^{2}}{M/r^{2}}\right)}}\ \ or\ \ \mu\left(\sqrt{\frac{M/{r}^{2}}{M_{u}/{R_{u}}^{2}}}\right)=\frac{1}{{1+\sqrt{\frac{M_{u}/{R_{u}}^{2}}{M/r^{2}}}}}$ (4)
@@
in which for most cases where local potentials are small, $\left(\frac{M_{u}}{R_{u}}+\frac{M}{r}\right)\sim\left(\frac{M_{u}}{R_{u}}\right)$. Inertial mass is then defined as the scalar $m_{i}=m_{g}G(M_{u}/R_{u})/c^{2}$ having an origin in the gravitation of all other masses in the universe, with $m_{g}$ the gravitational mass.
@@
It is trivial that, since only equivalent substitutions for $G$ and $a_{0}$ according to Sciama’s interpretation of Mach’s principle which yield around the same values are done (omitting integers and mathematical constants), and $M$ and $r$ take the same meaning as in Milgrom’s MOND, the resulting correction (3) and (4) for galaxy rotation curves are equivalent to Milgrom’s MOND. For $\frac{M_{u}/{R_{u}}^{2}}{M/r^{2}}<<1$, $\mu\left(\sqrt{\frac{M/{r}^{2}}{M_{u}/{R_{u}}^{2}}}\right)\rightarrow 1$ and equivalent Newtonian laws are restored with $v=c\sqrt{({M/r})/({M_{u}/R_{u}})}$. For $\frac{M_{u}/{R_{u}}^{2}}{M/r^{2}}>>1$, $\mu\left(\sqrt{\frac{M/{r}^{2}}{M_{u}/{R_{u}}^{2}}}\right)\rightarrow\sqrt{\frac{M/r^{2}}{M_{u}/{R_{u}}^{2}}}$, $v=c\sqrt[4]{M/M_{u}}$, and the speed of rotating stars around galaxies far away from their galactic centers is determined by the relative active gravitational masses between the local and the global system.
It is also trivial that integrating an average mass density of the homogeneous and isotropic universe at large enough scales over the causally connected spherical volume results in $NM_{u}/{R_{u}}^{2}$ (with $N$ being a dimensionless factor that can be omitted for simplicity, as it does not affect the orders of magnitude of the relationship between $M_{u}$ and $R_{u}$). A precise calculation would require considering Lienard-Wiechert potentials and the retarded Green functions to account for the propagation of gravity at the speed of light and current mass density instead of simple mass density. Under Machian MOND, the denominator $M_{u}/{R_{u}}^{2}$ in (4) is the total field intensity (the gravitational field intensity without the gravitational constant $G$) of all masses of the universe (omitting dimensionless factors for simplicity), taking the same physical meaning as $M/{r}^{2}$ in the numerator, which is the field intensity of the masses of the local system at the point in which the correction is applied.
@@
Thus, a transformation (4) can be identified between the non-inertial reference frame of a local system with a mass parameter $M$ and distance $r$, and a global reference frame of observable mass $M_{u}$ and size $R_{u}$ of the universe. The variable changed by this transformation is either the gravitational force (equivalent to a variable gravitational constant) or the inertial law. Note that in (1), the interpolating function in modified inertia depends on true acceleration, while in Machian MOND it depends on the local field intensity, which suggests that inertia has an origin in gravitation, as the modified inertia theories of Sciama and Treder suggest through the same parameters $M_{u}$ and $R_{u}$ present in Machian MOND.
@@
By decreasing the mass of the universe within the observable radius, the transformation (4) decreases the gravitational force when $\frac{M_{u}/{R_{u}}^{2}}{M/r^{2}}>>1$. But, considering also Sciama’s relationship for the gravitational constant, the gravitational force increases (having the same observational consequence as considering Sciama’s relationship alone, the decrease of inertia), and local velocities increase. The formal limit where $a_{0}\rightarrow 0$ in which Milgrom’s MOND approaches Newton’s laws does not hold, since the mass and size of the universe (for which $a_{0}$ is effectively replaced) always exist. The mass of the universe can never be zero in the system under study, it can only be small as the very same mass of the system itself. Decreasing the mass of the universe until the absence of any background ($M_{u}\rightarrow M,R_{u}\rightarrow r$), the transformation (4) reduces to some simple number $K$ which can be set to unity with the appropriate choice of integers in the formulation of the Machian transformation and Sciama’s Mach’s principle interpretation (a simple way to obtain $K=1$ is considering Schrödinger’s original definition of $G$ and multiplying the term in brackets of the transformation by $3$), and velocities approach $v\rightarrow Kc$. This is already a Machian feature of Sciama’s relationship by which without any background, rotation is undefined, and tangential speed can take any value up to the limit of the speed of light. The proposed transformation respects this result and also respects the scale invariance in mass, length, and time that Sciama’s relationship introduces in the Newtonian regime (classical Newtonian gravity is scale-dependent). Milgrom’s MOND is only length and time scale-invariant when $a_{0}\rightarrow\infty$ fixing $Ga_{0}$ to achieve constant velocities independent of radius at the flat part of the galaxy rotation curves. Machian MOND is thus scale-invariant in mass, length, and time $(m,r,t)\rightarrow(\lambda m,\lambda r,\lambda t)$ at all regimes. Doubling the mass or size of both the universe and the local system always results in the same estimated velocities.
@@ 3 Discussion
The relationship pointed out by Milgrom between his MOND’s acceleration constant $a_{0}$ and the cosmological constant or the Hubble parameter is interpreted merely as accidental due to the cosmological coincidence by which, at the present epoch, the energy densities associated with dark energy (the cosmological constant) and visible matter are of the same orders of magnitude.
@@
The Machian MOND’s transformation is a function $\mu(M_{u},R_{u},M,r)$ (4) between different scales and non-inertial reference frames: the relationship between a local frame in spacetime, and that of a global frame which is at rest (the background of the rest of the universe) with respect to the first one, with both being accelerated relative to each other. It expresses how gravity or inertia changes locally due to the background universe through relative masses and distances in both systems, in accordance with relational mechanics, and respects the weak equivalence principle since it does not depend on the inertial mass. The speed of light in Sciama’s relationship takes the meaning of the speed of gravity and causality justifying the choice of $R_{u}$ as the radius of the observable universe containing mass $M_{u}$, which is the part of the universe causally affecting local dynamics, for Mach’s interpretation not to violate locality.
@@
Although the origin of the Machian MOND transformation is not explained or derived here, it seems that it follows from Mach’s principle’s argument about which frame is acceleration defined with respect to. When the local field intensity is above that of the rest of the universe (at the cores of galaxies or within the solar system), the ‘main frame’ acceleration is defined with respect to is the local one, and one can describe the dynamics of the system without referring to the field intensity of the universe at large through Newtonian laws. But, if the field intensity of the rest of the universe is greater than the one from the local system (in the deep MOND regime, such as the flat velocity part of the disk in galaxies), it seems that acceleration is defined with respect to the background frame of the universe, and the field intensity of the universe at large must be taken into account non-linearly together with that of the local frame, to explain the dynamics of the local system (in the case of Machian MOND, effectively through the Machian MOND transformation (4) ). In between both regimes, the identification of the ‘main frame’ is shared between the local and global frames, described approximately by the slope of the Machian MOND transformation.
@@
Machian MOND satisfies several definitions of Mach’s principle: Newton’s gravitational constant $G$ is a dynamical field, local inertial frames are affected by the cosmic distribution of matter, and rotation is undefined up to the speed of light in the absence of the rest of the universe. If the correction applies to inertia instead of just gravity, then inertial mass is also affected by the global distribution of matter, and an isolated body in otherwise empty space would have no inertia. In agreement with Mach’s principle, if inertial mass is defined by the potential of the mass distribution (following Sciama’s argument), it is reasonable to think that it could also be affected by the field intensity of the mass distribution. Thus, Machian MOND introduces an even stronger (and non-linear, due to the external field effect of MOND) dependence of local dynamics on the distribution of matter than linear theories that just consider Sciama’s relationship, and it can be considered more Machian.
@@
Milgrom himself favors modifying inertia (which is the true aim of Mach’s principle), and not just gravity. An example of modified inertia is that of SR $F=m_{i}d(\gamma v)/dt$ with $\gamma$ being the Lorentz factor [34]. The original Milgrom’s MOND as modified inertia with the inertial term of the equation of motion $\mu\left(g/a_{0}\right)m_{i}g$ is translation, rotation, and Galilei invariant, but leads to non-conservation of momentum in a many-body systems because the equation of motion is not derivable from an action. As a modification of the kinetic term in the Lagrangian, it conserves momentum, energy, and angular momentum, but is not Galilei invariant. A non-relativistic, MOND, kinetic action for a particle with conservation laws of momentum, energy, and angular momentum, and the boost symmetry being Galilean, must be time non-local [10]. But this only occurs if it is a function of the trajectory $r(t)$ (since the function requires knowledge of the entire trajectory) and it is required to have a MOND limit when $a_{0}\rightarrow 0$, which is neither the case for Machian MOND. We have omitted the original negative sign of Sciama’s relationship $c^{2}\sim-GM_{u}/R_{u}$ for simplicity. But by considering it and changing the sign in Milgrom’s MOND standard interpolating function from positive to negative for consistency, the resulting transformation
@@
is equivalent to Machian MOND (4) and resembles the structure of a Lorentz-like transformation $\gamma=1/\sqrt{1-\beta}$. Another possibility is that this negative sign comes from $M/r^{2}$ having an opposite direction to that of $M_{u}/R_{u}^{2}$. If a function depending on velocities, such as the Lorentz factor, arises from Lorentz invariance in SR, where accelerations are absolute for consistency of the theory, one would expect a function depending on accelerations or field intensities, such as Machian MOND, to arise when relativizing accelerations. Lorentz invariance implies that the speed of light is constant in all inertial frames, and it seems that Machian MOND implies that the total field intensity at any point in the universe is always $>M_{u}/R_{u}^{2}$. It may be that Machian MOND arises from a new boost symmetry generalizing Lorentz invariance, for instance, through the relativity of acceleration, which is broken by the presence of the universe at large. Without a background universe, this symmetry would reduce to the Galilean or Lorentz one. If inertia has an origin in gravitation, then the modification on the kinetic term can be based only on the field intensity, without modifying the Poisson equation (and leaving gravity intact), which is shown in Machian MOND as modified inertia. The resulting Machian MOND formulation can be thought of as an in-between interpretation of modified inertia and modified gravity.
@@
Machian MOND, or a similar and equivalent approximate formulation, is considered to be a necessary non-linear feature that a phenomenological theory of modified inertia or modified gravity which incorporates Mach’s principle reduces to as an approximation, in order to agree with galaxy rotation dynamics. This might arise by imposing Mach’s principle (for instance, through the relativity of accelerations and inertia of local objects being determined by the sources of the gravitational field at the global scale) to a non-Machian non-linear Lagrangian or to a theory of gravity such as GR (in which solutions to local dynamics do not depend on the mass or size of the observable universe). In fact, analogous interpolating functions in SR and GR have, as independent variables, dimensionless ratios of the form $v/c$ and $MG/Rc^{2}$, similar to Machian MOND. The transformation function should arise naturally as an approximation, without the need of introducing it by hand (as in AQUAL or QUMOND), but the global $M_{u}$ and $R_{u}$ terms governing local dynamics should appear in the complete FundaMOND theory, which is not the case in GR. So far, only linear and non-Lorentz invariant Machian theories of modified inertia with Sciama’s relationship have been constructed, such as that of Hans-Jürgen Treder [21].

### v4

@@ 1.2 Mach’s Principle and modified inertia
Sciama’s relationship with today’s measurements of the current cosmological model is satisfied only in terms of orders of magnitude, because the Hubble tension impedes a precise calculation. It yields the orders of magnitude of the observed gravitational constant when considering not the baryonic and dark matter content for $M_{u}$, but the total energy content (since all forms of energy must gravitate according to GR), which roughly corresponds to the critical density of the universe as measured to be almost flat. As Milgrom stated in one of his conferences, “the only system that is strongly general relativistic and in the MOND regime is the Universe at large”. However, the critical density is calculated with the assumption of a constant gravitational constant itself and considering the dark matter content, which could be reduced through an effective correction to Newtonian laws, as will be proposed in the next section.
@@
Alexander Unzicker explored galaxy dynamics in relation to Mach’s Principle and the rotating bucket of water experiment, and proposed a basis for a MOND to solve flat rotation curves without considering Milgrom’s MOND [29]. He proposes the following thought experiment. According to Mach, the system of two distant masses rotating around their center of mass in equilibrium of gravitational and centrifugal forces in an empty universe, since rotation between them is undefined due to absence of absolute space, must be equivalent to the system of those two distant masses not rotating and without gravity (only radial or relative velocities are measurable, and tangential velocities are impossible to measure instantaneously, as they are indistinguishable from a rotation of the coordinate system of the observer). The maximum possible angular velocity is limited by $w_{max}=c/r$ so that the maximum tangential speed is $v=c$ for one of the masses rotating around the other one, since the speed of light cannot be surpassed in any case. This luminal rotational speed in absence of a background universe from Unzicker’s thought experiment is a feature of introducing Sciama’s relationship in Newton’s law of gravity, in which the Newtonian $v=\sqrt{GM/r}$ changes to $v=c\sqrt{(M/r)/(M_{u}/R_{u})}$ after introducing $G=c^{2}/(M_{u}/R_{u})$ and without a background universe $M_{u}\rightarrow M$ and $R_{u}\rightarrow r$, velocities are $v=c$. In other words, Sciama’s relationship has the equivalent effect of decreasing inertia when decreasing $M_{u}$.
@@
A complete FundaMOND theory remains only a faint hope until the origin of Milgrom’s MOND is correctly identified. Even though interesting ideas have been put forward relating MOND to Mach’s principle, an exact equivalence based on the Machian consequence that the inertia of a body should decrease when masses are removed from its neighborhood (at the deep MOND regime) will be presented, which should arise as an approximation in any non-linear theory of modified inertia or gravity which attempts to explain galaxy rotation curves. The motivation behind this interpretation of Milgrom’s MOND is that MOND can be reformulated with the same parameters of the mass and radius of the observable and causally connected universe that the Machian models of modified inertia contain.
@@ 2 Milgrom’s MOND from Mach’s Principle
Starting from the well-known form of Milgrom’s MOND standard or simple interpolating function correction to Newtonian gravity or to inertia, the ‘true’ gravitational acceleration $g$ or ‘true’ second law of inertia is
$g=\frac{G}{\mu\left(\frac{g}{a_{0}}\right)}\frac{M}{r^{2}}=\frac{g_{N}}{\mu\left(\frac{g}{a_{0}}\right)}\ \ or\ \ F=\mu\left(\frac{g}{a_{0}}\right)m_{i}g$ (1)
@@
with $a_{0}$ being Milgrom’s MOND acceleration scale constant, $g_{N}=GM/r^{2}$ the Newtonian gravitational acceleration, $M$ the active gravitational mass, $m_{i}$ the inertial mass, and $r$ the distance between the centers of mass of the active and passive gravitational masses. For the motion of a test particle in a gravitational field in the low acceleration or deep MOND regime, $\mu\left(a/a_{0}\right)\rightarrow a/a_{0}$ and $g=\sqrt{a_{0}g_{N}}$, so that with $g=v^{2}/r$, $v=\sqrt[4]{GMa_{0}}$ for achieving flat galaxy rotation curves in agreement with the Tully-Fisher law in a remarkable simple way with a single new free parameter. For the high acceleration regime, $\mu\left(a/a_{0}\right)\rightarrow 1$ and Newtonian laws are recovered.
Milgrom’s MOND acceleration scale constant $a_{0}$ is usually interpreted as a fundamental constant in MOND theory, but possibly related to the cosmological parameter of the cosmological constant $a_{0}\sim c^{2}\sqrt{\Lambda}$ or the energy density of the vacuum $\rho_{vac}$ with $\Lambda=8\pi G\rho_{vac}/c^{2}$ according to GR, which coincides in orders of magnitude (a better match can be done with $a_{0}=c^{2}\sqrt{\Lambda/3}$, although integers and Pi are onwards omitted for simplicity). This is due to Milgrom’s wish to define absolute acceleration with respect to a smooth vacuum energy frame in MOND as modified inertia. But Mach’s principle implies that acceleration should be defined with respect to the rest of the masses of the universe. Milgrom’s reluctance to explore a Machian version could be attributed to the widespread false claim that Mach’s principle necessarily implies anisotropy of inertia (due to non-uniform mass distributions), strongly constrained by the Hughes-Drever experiments. As shown by Treder, as long as inertial mass is a scalar, it exhibits no preferred direction, and no anisotropy of inertia. In order to express $a_{0}$ in terms of the mass and radius of the universe, $\rho_{vac}$ is substituted for the matter density of the universe $\rho_{vac}\sim\rho_{u}\sim M_{u}/{R_{u}}^{3}$ with $M_{u}$ the causally connected mass to the local system of study through the radius $R_{u}$ of the observable universe (or particle horizon radius), a relationship which also coincides in orders of magnitude. Newton’s gravitational constant $G$ is substituted for Sciama’s interpretation of Mach’s principle $G\sim c^{2}/(M_{u}/R_{u})$, so that $a_{0}\sim GM_{u}/{R_{u}}^{2}\sim c^{2}/R_{u}$, and by reverse engineering, $v\sim c\sqrt[4]{M/M_{u}}$ and $g\sim c^{2}\sqrt{M/M_{u}}/r$ in the equivalent deep MOND regime. The interpolating function is now scale-invariant in mass and length, resulting in a Machian MOND formulation
@@
$\mu\left(\sqrt{\frac{M/{r}^{2}}{M_{u}/{R_{u}}^{2}}}\right)=\frac{1}{\sqrt{1+\frac{M_{u}/{R_{u}}^{2}}{M/r^{2}}}}\ \ or\ \ \mu\left(\sqrt{\frac{M/{r}^{2}}{M_{u}/{R_{u}}^{2}}}\right)=\frac{1}{{1+\sqrt{\frac{M_{u}/{R_{u}}^{2}}{M/r^{2}}}}}$ (4)
@@
in which for most cases where local potentials are small, $\left(\frac{M_{u}}{R_{u}}+\frac{M}{r}\right)\sim\left(\frac{M_{u}}{R_{u}}\right)$.
@@
It is trivial that, since only equivalent substitutions for $G$ and $a_{0}$ according to Sciama’s relationship based on Mach’s principle which yield around the same values are done (omitting integers and mathematical constants), and $M$ and $r$ take the same meaning as in Milgrom’s MOND, the resulting corrections (3) and (4) for galaxy rotation curves are equivalent to Milgrom’s MOND. For $\frac{M_{u}/{R_{u}}^{2}}{M/r^{2}}<<1$, $\mu\left(\sqrt{\frac{M/{r}^{2}}{M_{u}/{R_{u}}^{2}}}\right)\rightarrow 1$ and Newtonian laws are restored with $v=c\sqrt{({M/r})/({M_{u}/R_{u}})}$. For $\frac{M_{u}/{R_{u}}^{2}}{M/r^{2}}>>1$, $\mu\left(\sqrt{\frac{M/{r}^{2}}{M_{u}/{R_{u}}^{2}}}\right)\rightarrow\sqrt{\frac{M/r^{2}}{M_{u}/{R_{u}}^{2}}}$, $v=c\sqrt[4]{M/M_{u}}$, and the speed of rotating stars around galaxies far away from their galactic centers is determined by the relative active gravitational masses between the local and the global system.
It is also trivial that the denominator $M_{u}/{R_{u}}^{2}$ in (4) is the total field intensity (the gravitational field intensity without the gravitational constant $G$) of all masses of the universe (omitting dimensionless factors for simplicity), taking the same physical meaning as $M/{r}^{2}$ in the numerator, which is the field intensity of the masses of the local system at the point in which the correction is applied.
@@
Thus, a transformation (4) can be identified between the non-inertial reference frame of a local system with a mass parameter $M$ and distance $r$, and a global reference frame of observable mass $M_{u}$ and size $R_{u}$ of the universe. The variable changed by this transformation is either the gravitational force (equivalent to a variable gravitational constant) or the inertia of a body. Note that in (1), the interpolating function depends on true acceleration, while in Machian MOND it depends on field intensity. In particular, as modified inertia, it depends on the inertial mass, the local mass, and the mass of the universe, showing that inertia arises from the overall mass distribution à la Mach. It also makes use of the same parameters $M_{u}$ and $R_{u}$ as the modified inertia theories of Sciama and Treder.
@@
By decreasing the mass of the universe within the observable radius, the transformation (4) decreases the gravitational force (or increases inertia) when $\frac{M_{u}/{R_{u}}^{2}}{M/r^{2}}>>1$. But, considering also Sciama’s relationship for the gravitational constant, the gravitational force increases (having the same observational consequence as considering Sciama’s relationship alone, the decrease of inertia), and local velocities increase. The formal limit where $a_{0}\rightarrow 0$ in which Milgrom’s MOND approaches Newton’s laws does not hold, since the mass and size of the universe (for which $a_{0}$ is effectively replaced) always exist. The mass of the universe can never be zero in the system under study, it can only be small as the very same mass of the system itself. Decreasing the mass of the universe until the absence of any background ($M_{u}\rightarrow M,R_{u}\rightarrow r$), the transformation (4) reduces to some simple number $K$ which can be set to unity with the appropriate choice of integers in the formulation of the Machian transformation and in Sciama’s relationship (a simple way to obtain $K=1$ is considering Schrödinger’s original definition of $G$ and multiplying the relational field term of the transformation by $3$), and velocities approach $v\rightarrow Kc$. This is already a Machian feature of Sciama’s relationship by which in absence of a background universe, rotation is undefined, and tangential speed can take any value up to the limit of the speed of light. The proposed transformation respects this result and also respects the scale invariance in mass, length, and time that Sciama’s relationship introduces in the Newtonian regime, while classical Newtonian gravity is scale-dependent. Milgrom’s MOND is only length and time scale-invariant when $a_{0}\rightarrow\infty$ fixing $Ga_{0}$ to achieve constant velocities independent of radius at the flat part of the galaxy rotation curves. Machian MOND is thus scale-invariant in mass, length, and time $(m,r,t)\rightarrow(\lambda m,\lambda r,\lambda t)$ at all regimes. Scaling the mass or size of both the universe and the local system always results in the same observational velocities.
@@ 3 Discussion
The relationship pointed out by Milgrom between his MOND’s acceleration scale constant $a_{0}$ and the cosmological constant or the Hubble parameter is interpreted merely as accidental due to the cosmological coincidence by which, at the present epoch, the energy densities associated with dark energy (the cosmological constant) and visible matter are of the same orders of magnitude.
@@
The Machian MOND’s transformation is a function $\mu(M_{u},R_{u},M,r)$ (4) between different scales and non-inertial reference frames: the relationship between a local frame in spacetime and that of a global frame which is at rest (the background of the rest of the universe) with respect to the first one, with both being accelerated relative to each other. It expresses how gravity or inertia changes locally due to the background universe through relative masses and distances in both systems, in accordance with relational mechanics, and respects the weak equivalence principle since it does not depend on the inertial mass. The speed of light in Sciama’s relationship takes the meaning of the speed of gravity and causality justifying the choice of $R_{u}$ as the radius of the observable universe containing mass $M_{u}$, which is the part of the universe causally affecting local dynamics, for Mach’s interpretation not to violate locality.
In Milgrom’s MOND, the external field effect, which is behind the non-linearity of MOND, does not allow to ignore the external field (the field of the mother system) when considering internal dynamics. But somehow the field of the universe, in which all systems and embedded into, is ignored in Milgrom’s MOND (if considered, all bodies in the universe would be above $GM_{u}/R_{u}^{2}$ and thus, in the high acceleration regime). By effectively replacing $a_{0}$ for $M_{u}/R_{u}^{2}$ in Machian MOND, the accelerated motion of the universe is considered due to the relativity of acceleration and inertia, when acceleration is defined with respect to the inertial frame of the universe à la Mach.
@@
Although the origin of the Machian MOND transformation is not explained or derived here, it seems that it follows from Mach’s principle’s argument about which frame is acceleration defined with respect to. When the local field intensity is above that of the rest of the universe (at the cores of galaxies or within the solar system), the ‘main frame’ acceleration is defined with respect to is the local one, and one can describe the dynamics of the system without referring to the field intensity of the universe at large through the usual Newtonian laws. But, if the field intensity of the rest of the universe is greater than the one from the local system (in the deep MOND regime, such as the flat velocity part of the disk in galaxies), it seems that acceleration is defined with respect to the background frame of the universe, and the field intensity of the universe at large must be taken into account non-linearly together with that of the local frame, to explain the dynamics of the local system (in the case of Machian MOND, effectively through the Machian MOND transformation (4) ). In between both regimes, the identification of the ‘main frame’ is shared between the local and global frames, described approximately by the slope of the Machian MOND transformation. This suggests that wherever local field intensities are higher than those of the universe at large, the Machian effects vanish (be that the one from Machian MOND, or from a varying gravitational constant in Brans-Dicke theory, in which tests showing the theory to be almost identical to GR are performed only at the high field intensity regime). And according to Mach, wherever local field intensities are below those of the universe at large, inertia is smaller and velocities are higher. This points towards the long-discussed problem of boundary conditions of the universe at large in GR, which has been linked to Mach’s principle in the past.
@@
Machian MOND satisfies several definitions of Mach’s principle: Newton’s gravitational constant $G$ is a dynamical field (and the inertial mass of a body increases with the agglomeration of masses in its neighborhood), local inertial frames are affected by the cosmic distribution of matter, and rotation is undefined up to the speed of light in the absence of the rest of the universe (or equivalently, a body in an empty universe has no inertia). In agreement with Mach’s principle, if inertial mass is defined by the potential of the mass distribution (following Sciama’s argument), it is reasonable to think that it could also be affected by the field intensity of the mass distribution. Thus, Machian MOND introduces an even stronger (and non-linear, due to the external field effect of MOND) dependence of local dynamics on the distribution of matter than linear theories that just consider Sciama’s relationship, and it can be considered more Machian.
@@
Milgrom himself favors modifying inertia (which is the true aim of Mach’s principle), and not just gravity. An example of modified inertia is that of SR $F=m_{i}d(\gamma v)/dt$ with $\gamma$ being the Lorentz factor [37]. The original Milgrom’s MOND as modified inertia with the inertial term of the equation of motion $\mu\left(g/a_{0}\right)m_{i}g$ is translation, rotation, and Galilei invariant, but leads to non-conservation of momentum in a many-body systems because the equation of motion is not derivable from an action. As a modification of the kinetic term in the Lagrangian, it conserves momentum, energy, and angular momentum, but is not Galilei invariant. A non-relativistic MONDian kinetic action for a particle with conservation laws of momentum, energy, and angular momentum, and the boost symmetry being Galilean, must be time non-local [10]. But this only occurs if it is a function of the trajectory $r(t)$ (since the function requires knowledge of the entire trajectory) and it is required to have a MOND limit when $a_{0}\rightarrow 0$, which is neither the case in Machian MOND. We have omitted the original negative sign of Sciama’s relationship $c^{2}\sim-GM_{u}/R_{u}$ for simplicity. But by considering it and changing the sign in the Machian MOND transformation (4) from positive to negative for consistency, the resulting transformation
@@
resembles the structure of a Lorentz-like transformation $\gamma=1/\sqrt{1-\beta}$. If a function depending on velocities, such as the Lorentz factor, arises from Lorentz invariance in SR, where accelerations are absolute for consistency of the theory, one would expect a function depending on accelerations or field intensities, such as Machian MOND, to arise when relativizing accelerations. Lorentz invariance implies that the speed of light is constant in all inertial frames, and it seems that Machian MOND implies that the total field intensity at any point in the universe is always $>M_{u}/R_{u}^{2}$. It may be that Machian MOND arises from a new boost symmetry generalizing Lorentz invariance, for instance, through the relativity of acceleration, which is broken by the presence of the universe at large. Without a background universe, this symmetry would reduce to the Galilean or Lorentzian one. If inertia has an origin in gravitation, then the modification on the kinetic term can be based only on field intensities, without modifying the Poisson equation (and leaving gravity intact), which is shown in Machian MOND as modified inertia. The resulting Machian MOND formulation can be thought of as an in-between interpretation of modified inertia and modified gravity.
@@
Machian MOND, or a similar and equivalent approximate formulation, is considered to be a necessary non-linear feature that a phenomenological theory of modified inertia or modified gravity which incorporates Mach’s principle reduces to as an approximation, in order to agree with galaxy rotation dynamics. This might arise by imposing Mach’s principle (for instance, through the relativity of accelerations and inertia of local objects being determined by the sources of the gravitational field at the global scale) to a non-Machian non-linear Lagrangian or to a theory of gravity such as GR (in which solutions to local dynamics do not depend on the mass or size of the observable universe). In fact, analogous interpolating functions in SR and GR have, as independent variables, dimensionless ratios of the form $v/c$ and $MG/Rc^{2}$, similar to Machian MOND. The transformation function should arise naturally as an approximation, without the need of introducing it by hand (as in AQUAL or QUMOND), but the global $M_{u}$ and $R_{u}$ parameters governing local dynamics should appear in the complete FundaMOND theory, which is not the case in GR. So far, only linear and non-Lorentz invariant Machian theories of modified inertia with Sciama’s relationship have been constructed, such as those of Treder and Schrödinger.